# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ 6.2.2. Gradient estimate via controllability and integration by parts
where in $(i)$, we applied the chain rule for the Malliavin derivative (see [61, Proposition 1.2.4]), and in $(ii)$, we denote by $\delta$ the adjoint operator of $\mathscr{D}$ on $L^{2}(\Omega;\mathbb{H}^{\sigma})$, usually referred to as the Skorohod operator, with its domain denoted by $\mathsf{D}(\delta)$. Since
@@
it follows that in the case (6.16) holds, the proof of Proposition 6.2 reduces to prove $g\in\mathsf{D}(\delta)$ and estimating $\mathbf{E}|\delta(g)|$. The latter will be performed using the following result, which can be proven as in [61, Propositions 1.3.1 and 1.3.11]. Below, we denote the subspace of progressively measurable processes with the subscript ‘${{\rm prog}}$’.
@@
As we will see in the proof of Proposition 5.6 below, the possibility of choosing $\sigma_{0}>\sigma+\frac{3}{2}$is crucial for accommodating diagonal operators, as demonstrated in Example 3.4. While this may not be surprising, it is important to emphasize that reaching the regime $\sigma_{0}\in(\sigma+1,\sigma+2)$ cannot be done directly and requires a two-step argument. Moreover, the regularity threshold is optimal because $Q\in\mathscr{L}_{2}(\mathbb{L}^{2},\mathbb{H}^{\sigma})$, and therefore the velocity $v$ belongs to $\mathbb{H}^{1+\sigma}$ at most in space.
@@
To prove the bounds, here we apply stochastic maximal $L^{p}$-regularity estimates. To this end, fix $\sigma_{1}<\sigma_{0}$, $p_{0}\in(2,\infty)$ and $\kappa_{0}\in[0,\frac{p_{0}}{2}-1)$ such that $\sigma_{0}<\sigma_{1}+1-\frac{2}{p_{0}}$ and $\sigma\geq\sigma_{1}+1-2\frac{1+\kappa_{0}}{p_{0}}$. Now, by stochastic maximal $L^{p}$-regularity for the Laplacian applied to the ground space $H^{\sigma_{1}-1}(\mathbb{T}^{3};{\mathbb{R}}^{2})$ with time weight $\kappa_{0}$ (see e.g., [7, Theorem 7.16] and the periodic version of [48, Example 10.1.5]), we obtain
@@
Step 2: Conclusion. The idea here is to use again maximal $L^{2}$-regularity estimates working on $[T/2,T]$ instead of $[0,T]$, and exploiting the estimates of Step 1. More precisely, for all $\sigma_{0}\in(\sigma+1,\sigma+2)$, maximal $L^{2}$-regularity estimates for the Laplace operator on $H^{\sigma_{0}-1}(\mathbb{T}^{3})$ (see e.g., [66, Theorem 3.5.7]) ensures the existence of a constant $M>0$ independent of $T,v$ and $h$ such that
@@
There exists $r_{0}$ depending only on the parameters introduced in Subsection 2.2.2 for which the following holds. For all $r>r_{0}$, there exist constants $a_{0},b_{0},T_{0}>0$ such that for all $0<T\leq T_{0}$, $V\in\mathcal{H}$ and $h\in T_{V}\mathcal{H}$ there exists a control $g_{t}\in\mathsf{D}(\mathscr{D})$ for a.a. $t\in(0,T)$ and $(\mathscr{D}_{s}g_{t})_{s,t}:\Omega\times(0,T)^{2}\to\mathscr{L}_{2}(\mathbb{L}^{2}_{0})$ has a measurable modification, and
@@ 6.4.1. Construction of the control
To prove Lemma 6.6, we follow the splitting in high and low frequencies as in [16, 29, 67]. The idea is to split the system (6.5) into a high-frequency part where the noise is active in all components and a low-order part that is finite dimensional and only a few modes are active but they are still sufficient to obtain improvement on the smoothness via an Hörmander type condition, see Lemma 6.10 and Remark 6.12.
@@
Let us remark that the SPDEs in (6.32) are coupled via the nonlinearities $F_{\mathrm{L}}(V)$ and $F_{\mathrm{H}}(V)$. Moreover,
@@
(6.34) $\mathcal{H}_{\mathrm{L}}\stackrel{{\scriptstyle{\rm def}}}{{=}}(\Pi_{\mathrm{L}}\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))\times{\mathbb{R}}^{3}\times\mathbb{S}^{2}\times{\mathbb{R}}^{6}\qquad\text{ and }\qquad\mathcal{H}_{\mathrm{H}}\stackrel{{\scriptstyle{\rm def}}}{{=}}\Pi_{\mathrm{H}}\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}),$
@@
Note that $R_{t}^{\mathrm{L}}$ and $R_{t}^{\mathrm{H}}$ are not the projection of the Jacobian $J_{t}$ onto the spaces $\mathcal{H}_{\mathrm{L}}$ and $\mathcal{H}_{\mathrm{H}}$, respectively. Indeed, the projections $\Pi_{\mathrm{L}}$ and $\Pi_{\mathrm{H}}$ do not commute with the operator $J\mapsto DF(v)J$. However, as we expect $J_{t}\approx\mathrm{Id}$ for small times, then $\Pi_{\mathrm{L}}R_{t}^{\mathrm{H}}\approx 0$ and $\Pi_{\mathrm{H}}R_{t}^{\mathrm{L}}\approx 0$ for small times and therefore $R^{\mathrm{H}}_{t}$ and ${\mathbb{R}}_{t}^{\mathrm{L}}$ are good approximation of the full Jacobian $J_{t}$.
For notational convience, for $V^{\mathrm{L}}\in\mathcal{H}_{\mathrm{L}}$ and $t\geq 0$, we let
$T_{V^{\rm L}_{t}}\mathcal{H}_{\mathrm{L}}=\Pi_{{\rm L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))\times{\mathbb{R}}^{3}\times T_{\xi(t)}\mathbb{S}^{2}\times{\mathbb{R}}^{6}.$
where $V_{t}^{\mathrm{L}}$ is given by the solution to the SPDE (6.32).
From the above definitions, for all $s,t\geq 0$, we have $R_{t,s}$ is a linear mapping from $T_{V_{s}^{\rm L}}\mathcal{H}_{\mathrm{L}}$ to $T_{V_{t}^{\rm L}}\mathcal{H}_{\mathrm{L}}$, and $Q_{\mathrm{L}}:T_{V^{\mathrm{L}}_{s}}\mathcal{H}\to T_{V^{\mathrm{L}}_{t}}\mathcal{H}$.
Finally, we also need the inverse of the operator $R^{\mathrm{L}}_{s,t}$. Let
@@
(6.37) $\partial_{t}S_{s,t}^{\mathrm{L}}=S_{s,t}^{\mathrm{L}}A_{\mathrm{L}}+S_{t,s}^{\mathrm{L}}D_{\mathrm{L}}F(V).$
Next, let us comment on the well-definiteness of the inverse in (6.36). Note that $R^{\mathrm{L}}_{s,t}$ is a-priori invertible only for short times as $R^{\mathrm{L}}_{s,t}\to\mathrm{Id}$ as $t\downarrow s$. However, the global well-posedness of the problem (6.37) shows the global in-time invertibility of the approximate Jacobian $R^{\mathrm{L}}_{s,t}$.
@@
$\mathbf{E}|(\mathscr{M}^{{\rm L}}_{T})^{-1}|^{p}\lesssim_{p}T^{-ap}(1+|\eta|)^{bp}\ \text{ for all }p>1.$
@@
As for the high-mode part, we would like to choose a control similar to the one used in (6.22), where, ideally, $(Q,\widetilde{J})$ are replaced by $(Q^{\mathrm{H}},R^{\mathrm{H}})$. However, in the current situation, it turns out that the error of the low frequencies to the high ones is too large. This issue arises because the parameter $a>0$ in Lemma 6.7 can be quite large, making the control of the partial Mallavin matrix, which governs the low frequencies, so singular as $T\downarrow 0$ that the smoothing effect of $A_{\mathrm{H}}$ cannot compensate its contribution. To solve this issue, as in [16, Section 6], we use a frequency cutoff in the high-frequencies control:
@@
The system (6.44) is constructed in such a way that the natural candidate for the solutions $\gamma_{t}$ and $\delta_{t}$ are $\mathscr{D}_{g}V^{\mathrm{L}}_{t}$ and $\mathscr{D}_{g}V^{\mathrm{H}}_{t}$, respectively; cf., (6.40) for the first equation. Note that we could not directly define (6.43) with $\gamma_{t}$ replaced by $\mathscr{D}_{g}V^{\mathrm{L}}_{t}$, because defining the latter requires knowing the high part of the control $(g_{t})_{t}$, which we aim to define. Therefore, by considering the system (6.44), we avoid the previous circular argument. The usefulness of the first term in the high-frequency control (6.43) will be clear below when constructing the remainder $r_{T}=DV_{T}h-\mathscr{D}_{g}V_{T}$, cf., Lemma 6.6.
@@
In the above, $C([0,T];T_{V_{t}}\mathcal{H})$ stands for the set of maps continuous maps $f$ on $[0,T]$ such that $f_{t}\in T_{V_{t}}\mathcal{H}$ for all $t\in[0,T]$. The above result verbatim from the argument in [16, Lemma 6.10], where instead of the nonlinear estimates for the Navier-Stokes nonlinearity, one uses Lemma 5.2 and the truncation in $\mathbb{H}^{\sigma}$ of the nonlinearity in (6.4). Therefore, we omit the details.
@@
The proof of Lemma 6.7 the above result via the following
@@
Moreover, for all $\varepsilon>0$ and $V\in\mathcal{H}_{\mathrm{L}}$, there exist costants $a_{1},b_{1}>1$ such that
(6.48) $\sup_{|h|_{\mathcal{H}_{\mathrm{L}}}=1}\mathbf{P}\big{(}\langle\mathscr{M}^{{\rm L}}_{T}h,h\rangle<\varepsilon\big{)}\lesssim_{p}t^{-a_{1}p}(1+|\eta|)^{b_{1}p}\varepsilon^{p},$
@@
Let us first show how Lemma 6.9 implies Lemma 6.7. The argument below largely follows the one in [39, Lemma 4.7] where we keep track of the constants which are needed because of the explosion as $T\to 0$.
@@
Let $N=\dim\mathcal{H}^{\rm L}$. Arguing as in [39, Lemma 4.7], by a covering argument, there exists an absolute constant $C_{0}>0$ such that, for all $\varepsilon>0$,
@@
$\lesssim t^{-ap}(1+|\eta|)^{bp},$
@@
where $(a_{j}^{(\eta)})_{j=1}^{6}$ is the standard basis of ${\mathbb{R}}^{6}$. The zeros in the above definition ensure that $\mathbf{q}_{j}\in T_{V_{\mathrm{L}}}\mathcal{H}$ for all $V_{\mathrm{L}}\in\Pi_{\mathrm{L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))$.
@@
(6.51) $\langle\mathscr{M}^{{\rm L}}_{t}h,h\rangle=\sum_{j\in\mathcal{J}_{\mathrm{L}}}\int_{0}^{t}\langle S^{\mathrm{L}}_{s}\mathbf{q}^{j}h,h\rangle^{2}\,\mathrm{d}s$
@@
$X_{t}^{j}\stackrel{{\scriptstyle{\rm def}}}{{=}}\langle S_{t}^{\rm L}\mathbf{q}^{j},h\rangle,$
@@
Now, arguing as in [16, Proposition 6.12], the key point is to study Hörmander type conditions for the vectors $(\mathbf{q}^{j},A_{\mathrm{L}}\mathbf{q}^{j}-D_{\mathrm{L}}F_{\mathrm{L}}(V_{t})\mathbf{q}^{j})_{j\in\mathcal{J}_{\mathrm{L}}}$. The following is the key algebraic result to prove (6.48), cf., [16, Lemma 6.15].
@@
(6.52) $\max_{j\in\mathcal{J}_{\mathrm{L}}}\big{\{}|\langle\mathbf{q}^{j},h\rangle|,|\langle[A_{\mathrm{L}}-D_{\mathrm{L}}F_{\mathrm{L}}(v,\mathbf{x},\xi)]\mathbf{q}_{j},h\rangle|\big{\}}\geq\frac{C_{0}}{(1+|\eta|)^{5}}|h|$
@@
The implication from Lemma 6.10 to Lemma 6.9 is well-known and follows verbatim from the proof of [16, Lemma 6.15]. The details are omitted for brevity.
@@
$+\chi_{r}(\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})})H(\xi,\eta),$
where $F_{\mathrm{L}}^{(v)}$ and $F^{(p)}_{\mathrm{L}}$ is the velocity and projective part of the nonlinearity low mode projection of the nonlinearity $F$ in (6.2), i.e.,
@@
The main step in the proof of Lemma 6.10 is the following result.
@@
There exist positive constants $r_{0},c_{0}$ such that, for all $r\geq r_{0}$, $V=(v_{\mathrm{L}},\mathbf{x},\xi,\eta)\in\mathcal{H}$ and $h=(v_{\mathrm{L}}^{(h)},\mathbf{x}^{(h)},\xi^{(h)},\eta^{(h)})\in T_{V}\mathcal{H}_{\mathrm{L}}$,
(6.55) $\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}^{j},h\rangle|$ $\geq c_{0}\big{(}|v_{\mathrm{L}}^{(h)}|+|\eta^{(h)}|\big{)},$
(6.56) $\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j},h\rangle|$ $\geq c_{0}\big{(}|\mathbf{x}^{(h)}|+|\xi^{(h)}|\big{)}.$
Proof.
The claim (6.55) follows clearly from $\inf_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}q_{(\mathbf{k},\ell)}>0$ and (6.50). Next, we focus on (6.56). To this end, let us first collect some facts. If $v\in\Pi_{\mathrm{L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))$, then one has the decomposition:
(6.57) $v(\mathbf{x})=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}v_{(\mathbf{k},\ell)}\gamma_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x})$
where $\mathbf{x}\in\mathbb{T}^{3}$ and $v_{(\mathbf{k},\ell)}=\int_{\mathbb{T}^{3}}v(\mathbf{x})\cdot\gamma_{(\mathbf{k},\ell)}e_{(\mathbf{k},\ell)}(\mathbf{x})\,\mathrm{d}\mathbf{x}$. By (2.1), the third component $w$ of the corresponding velocity field $u=(v,w)$ is given by
@@
$w(\mathbf{x})=-\int_{0}^{z}\nabla_{x,y}\cdot v(x,y,z^{\prime})\,\mathrm{d}z^{\prime}=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}v_{(\mathbf{k},\ell)}\gamma^{(w)}_{(\mathbf{k},\ell)}e_{(\mathbf{k},\ell)}(\mathbf{x})$
for $\mathbf{x}=(x,y,z)\in\mathbb{T}^{3}$, where

### v2

@@ 6.2.2. Gradient estimate via controllability and integration by parts
where in $(i)$, we applied the chain rule for the Malliavin derivative (see [62, Proposition 1.2.4]), and in $(ii)$, we denote by $\delta$ the adjoint operator of $\mathscr{D}$ on $L^{2}(\Omega;\mathbb{H}^{\sigma})$, usually referred to as the Skorohod operator, with its domain denoted by $\mathsf{D}(\delta)$. Since
@@
it follows that in the case (6.16) holds, the proof of Proposition 6.2 reduces to prove $g\in\mathsf{D}(\delta)$ and estimating $\mathbf{E}|\delta(g)|$. The latter will be performed using the following result, which can be proven as in [62, Propositions 1.3.1 and 1.3.11]. Below, we denote the subspace of progressively measurable processes with the subscript ‘${{\rm prog}}$’.
@@
As we will see in the proof of Proposition 5.6 below, the possibility of choosing $\sigma_{0}>\sigma+\frac{3}{2}$is crucial for accommodating diagonal operators, as demonstrated in Example 3.4. While this may not be surprising, it is important to emphasise that reaching the regime $\sigma_{0}\in(\sigma+1,\sigma+2)$ cannot be done directly and requires a two-step argument. Moreover, the regularity threshold is optimal because $Q\in\mathscr{L}_{2}(\mathbb{L}^{2},\mathbb{H}^{\sigma})$, and therefore the velocity $v$ belongs to $\mathbb{H}^{1+\sigma}$ at most in space.
@@
To prove the bounds, here we apply stochastic maximal $L^{p}$-regularity estimates. To this end, fix $\sigma_{1}<\sigma_{0}$, $p_{0}\in(2,\infty)$ and $\kappa_{0}\in[0,\frac{p_{0}}{2}-1)$ such that $\sigma_{0}<\sigma_{1}+1-\frac{2}{p_{0}}$ and $\sigma\geq\sigma_{1}+1-2\frac{1+\kappa_{0}}{p_{0}}$. Now, by stochastic maximal $L^{p}$-regularity for the Laplacian on the ground space $H^{\sigma_{1}-1}(\mathbb{T}^{3};{\mathbb{R}}^{2})$ with time weight $\kappa_{0}$ (see e.g., [7, Theorem 7.16] and [48, Example 10.1.5]), we obtain
@@
Step 2: Conclusion. The idea here is to use again maximal $L^{2}$-regularity estimates working on $[T/2,T]$ instead of $[0,T]$, and exploiting the estimates of Step 1. More precisely, for all $\sigma_{0}\in(\sigma+1,\sigma+2)$, maximal $L^{2}$-regularity estimates for the Laplace operator on $H^{\sigma_{0}-1}(\mathbb{T}^{3})$ (see e.g., [67, Theorem 3.5.7]) ensures the existence of a constant $M>0$ independent of $T,v$ and $h$ such that
@@
There exists $r_{0}$ depending only on the parameters introduced in Subsection 2.2.2 for which the following holds. For all $r>r_{0}$, there exist constants $a_{0},b_{0},T_{0}>0$ such that for all $0<T\leq T_{0}$, $V\in\mathcal{H}$ and $h\in T_{V}\mathcal{H}$ there exists a control $(g_{t})_{t}\in\mathsf{D}(\mathscr{D})$ for a.a. $t\in(0,T)$ and $(\mathscr{D}_{s}g_{t})_{s,t}:\Omega\times(0,T)^{2}\to\mathscr{L}_{2}(\mathbb{L}^{2}_{0})$ has a measurable modification, and
@@ 6.4.1. Construction of the control
To prove Lemma 6.6, we follow the splitting in high and low frequencies as in [16, 29, 68]. The idea is to split the system (6.5) into a high-frequency part where the noise is active in all components and a low-order part that is finite dimensional and only a few modes are active but they are still sufficient to obtain improvement on the smoothness via an Hörmander type condition, see Lemma 6.10.
@@
Let us remark that the SPDE in (6.32) are coupled via the nonlinearities $F_{\mathrm{L}}(V)$ and $F_{\mathrm{H}}(V)$. Moreover,
@@
(6.34) $\mathcal{H}_{\mathrm{L}}\stackrel{{\scriptstyle{\rm def}}}{{=}}(\Pi_{\mathrm{L}}\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))\times\mathcal{O}\times\mathbb{S}^{2}\times{\mathbb{R}}^{6}\qquad\text{ and }\qquad\mathcal{H}_{\mathrm{H}}\stackrel{{\scriptstyle{\rm def}}}{{=}}\Pi_{\mathrm{H}}\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}),$
@@
Note that $R_{0,t}^{\mathrm{L}}$ and $R_{0,t}^{\mathrm{H}}$ are not the projection of the Jacobian $J_{0,t}=DV_{t}$ onto the spaces $\mathcal{H}_{\mathrm{L}}$ and $\mathcal{H}_{\mathrm{H}}$, respectively. Indeed, the projections $\Pi_{\mathrm{L}}$ and $\Pi_{\mathrm{H}}$ do not commute with the operator $J\mapsto DF(v)J$. However, as we expect $J_{0,t}\approx\mathrm{Id}$ for small times, then $\Pi_{\mathrm{L}}R_{0,t}^{\mathrm{H}}\approx 0$ and $\Pi_{\mathrm{H}}R_{0,t}^{\mathrm{L}}\approx 0$ for small times and therefore $R^{\mathrm{H}}_{0,t}$ and $R_{0,t}^{\mathrm{L}}$ are good approximation of the full Jacobian $J_{0,t}$.
For notational convience, for $V^{\mathrm{L}}\in\mathcal{H}_{\mathrm{L}}$, we let
$T_{V^{\rm L}}\mathcal{H}_{\mathrm{L}}=\Pi_{{\rm L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))\times{\mathbb{R}}^{3}\times T_{\xi}\mathbb{S}^{2}\times{\mathbb{R}}^{6}.$
From the above definitions, for all $s,t\geq 0$, we have $R_{t,s}$ is a linear mapping from $T_{V_{s}^{\rm L}}\mathcal{H}_{\mathrm{L}}$ to $T_{V_{t}^{\rm L}}\mathcal{H}_{\mathrm{L}}$, and $Q_{\mathrm{L}}:T_{V^{\mathrm{L}}_{s}}\mathcal{H}\to T_{V^{\mathrm{L}}_{t}}\mathcal{H}$. Finally, we also need the inverse of the operator $R^{\mathrm{L}}_{s,t}$. Let
@@
(6.37) $\partial_{t}S_{s,t}^{\mathrm{L}}=S_{s,t}^{\mathrm{L}}A_{\mathrm{L}}+S_{t,s}^{\mathrm{L}}D_{\mathrm{L}}F(V),\qquad S_{s,s}^{\mathrm{L}}=\mathrm{Id}.$
Next, let us comment on the well-definedness of the inverse in (6.36). Note that $R^{\mathrm{L}}_{s,t}$ is a-priori invertible only for short times as $R^{\mathrm{L}}_{s,t}\to\mathrm{Id}$ as $t\downarrow s$. However, the global well-posedness of the problem (6.37) shows the global in-time invertibility of the approximate Jacobian $R^{\mathrm{L}}_{s,t}$.
@@
$\mathbf{E}|(\mathscr{M}^{{\rm L}}_{T})^{-1}|^{p}\lesssim_{p}T^{-ap}\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})^{-b}(1+|\eta|)^{bp}\ \text{ for all }p>1.$
@@
As for the high-mode part, we would like to choose a control similar to the one used in (6.22), where, ideally, $(Q,\widetilde{J})$ are replaced by $(Q^{\mathrm{H}},R^{\mathrm{H}})$. However, in the current situation, it turns out that the error of the low frequencies compared to the high ones is too large. This issue arises because the parameter $a>0$ in Lemma 6.7 can be quite large, making the control of the partial Mallavin matrix, which governs the low frequencies, so singular as $T\downarrow 0$ that the smoothing effect of $A_{\mathrm{H}}$ cannot compensate its contribution. To solve this issue, as in [16, Section 6], we use a frequency cutoff in the high-frequencies control:
@@
The system (6.44) is constructed in such a way that the natural candidates for the solutions $\gamma_{t}$ and $\delta_{t}$ are $\mathscr{D}_{g}V^{\mathrm{L}}_{t}$ and $\mathscr{D}_{g}V^{\mathrm{H}}_{t}$, respectively; cf., (6.40) for the first equation. Note that we could not directly define (6.43) with $\gamma_{t}$ replaced by $\mathscr{D}_{g}V^{\mathrm{L}}_{t}$, because defining the latter requires knowing the high part of the control $(g_{t})_{t}$, which we aim to define. Therefore, by considering the system (6.44), we avoid the previous circular argument. The usefulness of the first term in the high-frequency control (6.43) will be clear below when constructing the remainder $r_{T}=DV_{T}h-\mathscr{D}_{g}V_{T}$, cf., Lemma 6.6.
@@
In the above, $C([0,T];T_{V_{t}}\mathcal{H})$ stands for the set of maps continuous maps $f$ on $[0,T]$ such that $f_{t}\in T_{V_{t}}\mathcal{H}$ for all $t\in[0,T]$. The above result is verbatim from the argument in [16, Lemma 6.10], where instead of the nonlinear estimates for the Navier-Stokes nonlinearity, one uses Lemma 5.2 and the truncation in $\mathbb{H}^{\sigma}$ of the nonlinearity in (6.4). Therefore, we omit the details.
@@
The proof of Lemma 6.7 of the above result via the following
@@
Moreover, there exist costants $a_{1},b_{1}>1$ such that, for all $\varepsilon>0$ and $V\in\mathcal{H}_{\mathrm{L}}$,
(6.48) $\sup_{|h|_{\mathcal{H}_{\mathrm{L}}}=1}\mathbf{P}\big{(}\langle\mathscr{M}^{{\rm L}}_{T}h,h\rangle<\varepsilon\big{)}\lesssim_{p}t^{-a_{1}p}\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})^{-b_{1}}(1+|\eta|)^{b_{1}p}\varepsilon^{p},$
@@
Let us first show how Lemma 6.9 implies Lemma 6.7. The argument below largely follows the one in [39, Lemma 4.7], where we keep track of the constants which are needed because of the explosion as $T\to 0$.
@@
Let $N=\dim\mathcal{H}_{\rm L}$, and recall that $T_{V}\mathcal{H}_{\mathrm{L}}=\Pi_{{\rm L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))\times{\mathbb{R}}^{3}\times T_{\xi}\mathbb{S}^{2}\times{\mathbb{R}}^{6}$ for $V\in\mathcal{H}_{\mathrm{L}}$. Arguing as in [39, Lemma 4.7], by a covering argument, there exists an absolute constant $C_{0}>0$ such that, for all $\varepsilon>0$,
@@
$\lesssim t^{-a_{1}(p+2N+1)}\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})^{-b_{1}(p+2N+1)}(1+|\eta|)^{b_{1}(p+2N+1)},$
@@
where $(a_{j}^{(\eta)})_{j=1}^{6}$ is the standard basis of ${\mathbb{R}}^{6}$. The zeros in the above definition ensure that $\mathbf{q}_{j}\in T_{V_{\mathrm{L}}}\mathcal{H}$ for all $V_{\mathrm{L}}\in\Pi_{\mathrm{L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))$. Note that $\mathcal{J}_{\mathrm{L}}$ gather all the ‘directions’ in the low-part of the system (6.32) where the noise is present.
@@
(6.51) $\langle\mathscr{M}^{{\rm L}}_{t}h,h\rangle=\sum_{j\in\mathcal{J}_{\mathrm{L}}}\int_{0}^{t}\langle S^{\mathrm{L}}_{s}\mathbf{q}_{j}h,h\rangle^{2}\,\mathrm{d}s$
@@
$X_{t}^{j}\stackrel{{\scriptstyle{\rm def}}}{{=}}\langle S_{0,t}^{\rm L}\mathbf{q}_{j},h\rangle,$
@@
Now, arguing as in [16, Proposition 6.12], the key point is to study Hörmander type conditions for the vectors $(\mathbf{q}_{j},A_{\mathrm{L}}\mathbf{q}_{j}-D_{\mathrm{L}}F_{\mathrm{L}}(V_{t})\mathbf{q}_{j})_{j\in\mathcal{J}_{\mathrm{L}}}$ appearing in the above. The following is the key algebraic result to prove (6.48), cf., [16, Lemma 6.15].
@@
(6.52) $\max_{j\in\mathcal{J}_{\mathrm{L}}}\big{\{}|\langle\mathbf{q}_{j},h\rangle|,|\langle[A_{\mathrm{L}}-D_{\mathrm{L}}F_{\mathrm{L}}(v,\mathbf{x},\xi)]\mathbf{q}_{j},h\rangle|\big{\}}\geq\frac{C_{0}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}}{(1+|\eta|)^{3}}|h|$
@@
The implication from Lemma 6.10 to Lemma 6.9 is well-known and follows verbatim from the proof of [16, Lemma 6.15]. The details are omitted for brevity. We note that the Lie brackets, which typically appear in algebraic lemmas of this kind, are absent in Lemma 6.10 because the functional $\langle\cdot,\mathbf{q}_{j}\rangle$ used to obtain the system for $X_{t}^{j}$ does not depend on the solution $(V_{t})_{t}$ of the truncated SPDE (6.5). As a result, the first-order commutators reduce to $[A_{\mathrm{L}}-D_{\mathrm{L}}F_{\mathrm{L}}(v,\mathbf{x},\xi)]\mathbf{q}_{j}$.
@@
$+\chi_{r}(\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})})H(\mathbf{x},\xi,\eta),$
where $F^{{\rm vel}}_{\mathrm{L}}$ and $F^{{\rm pr}}_{\mathrm{L}}$ are the velocity and projective part of the nonlinearity low mode projection of the nonlinearity $F$ in (6.2), i.e.,
@@
The following constitutes the main ingredient in the proof of Lemma 6.10.
@@
There exist positive constants $r_{0},c_{0}$ such that, for all $r\geq r_{0}$, $V=(v,\mathbf{x},\xi,\eta)\in\mathcal{H}$ and $h=(h_{v},h_{\mathbf{x}},h_{\xi},h_{\eta})\in T_{V}\mathcal{H}_{\mathrm{L}}$,
(6.55) $\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}_{j},h\rangle|$ $\geq c_{0}\big{(}|h_{v}|+|h_{\eta}|\big{)},$
(6.56) $\max_{j\in\mathcal{Z}_{\mathrm{L}}}|\langle D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j},h\rangle|$ $\geq c_{0}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}\big{(}|h_{\mathbf{x}}|+|h_{\xi}|\big{)}.$
We emphasise that in (6.56), the maximum is taken exclusively over $\mathcal{Z}_{\mathrm{L}}$, deliberately excluding the $\eta$-directions in $\mathcal{J}_{\mathrm{L}}\setminus\mathcal{Z}_{\mathrm{L}}$. This will be crucial later to prove (6.52) for low values of $\|v\|_{\mathbb{H}^{\sigma}_{0}}$ at which the truncation in (6.4) is not active.
Before going into the proof of Lemma 6.11, we collect some facts. If $v\in\Pi_{\mathrm{L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))$, then for some constants $(v_{(\mathbf{k},\ell)})_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}$ it holds that (see also the beginning of Subsection 2.2 for the notation)
(6.57) $v(\mathbf{x})=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}v_{(\mathbf{k},\ell)}\gamma_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x})\ \ \text{ for }\ \mathbf{x}\in\mathbb{T}^{3}.$
By (2.1), the third component of the corresponding velocity field is given by
@@
$\textstyle w(\mathbf{x})=-\int_{0}^{z}\nabla_{x,y}\cdot v(x,y,z^{\prime})\,\mathrm{d}z^{\prime}=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}\gamma^{(w)}_{(\mathbf{k},\ell)}v_{(\mathbf{k},\ell)}\big{(}e_{\mathbf{k}}(\mathbf{x})-e_{\mathbf{k}}(\mathbf{x}_{0})\big{)}$
[…]
$\mathbf{x}_{0}$ =$\stackrel{{\scriptstyle{\rm def}}}{{=}}(x,y,0),$